OpenStudy (anonymous):

can anyone help me to the answer for this... (will post equation in a sec...)

OpenStudy (anonymous):

\[\sqrt{\sqrt{x-2}+x}=2\]

OpenStudy (anonymous):

Square both sides

OpenStudy (anonymous):

root(x-2) + x =4 root(x-2) = 4-x square again

OpenStudy (anonymous):

x-2 = 16 - 8x + x^2 9x = 18 + x^2 x^2 - 9x + 18 = 0

OpenStudy (anonymous):

(x-6)(x-3)=0 x=6 or x=3

OpenStudy (anonymous):

@niravshah08 how on earth did you get the first line x-2 = 16 - 8x + x^2 ???

OpenStudy (anonymous):